OpenStudy (strawberry17):

Explain how to solve please She purchased a jet ski for $1,674 using a 12-month deferment plan. The interest rate is 23.37%. A down payment of $375 is required as well as a minimum monthly payment of $53. What is the balance at the beginning of the 13th month if only the minimum payment is made during the introductory period? $1,001.30 $1,098.96 $1,584.30 $1,637.30

OpenStudy (strawberry17):

I am solving it and not getting any one of these..

OpenStudy (strawberry17):

I'm guessing it could be the second or third one but not sure..

OpenStudy (anonymous):

it's the second one :P

OpenStudy (strawberry17):

how do you know?

OpenStudy (anonymous):

because ima smart person :D

OpenStudy (strawberry17):

are you for sure?? I really want to get this one right..

OpenStudy (strawberry17):

are u there?

OpenStudy (anonymous):

oh sorry yeah i'm sure :)

OpenStudy (strawberry17):

Do you know how to solve it, possibly?

OpenStudy (anonymous):

not really

OpenStudy (strawberry17):

@jperez570 can u help?

OpenStudy (strawberry17):

are u there?

OpenStudy (anonymous):

As soon as i'm done with this AP homework I will

OpenStudy (strawberry17):

k

OpenStudy (tkhunny):

Charging 0.2337% monthly and paying $53, I get $928.55 Charging annual interest on the remaining balance after 1 year, I get $817.94 Double counting interest in a way that amounts to Loan Sharking, I get $1,311.19 Having said that last one, I sincerely hope it's not the last two. You'll have to describe EXACTLY what happens to interest during the deferral period.

OpenStudy (strawberry17):

Well, since she did not pay off the whole balance during the deferment period, she would have to pay interest on the principal balance. So I did: $1674 • .02337 = 391.21 1674 + 391.21 = 2065.21 2065.21 - (53 • 12) = 1429.21 So, I got 1,429.21 and that is closest to the third one, if I even solved it right..

OpenStudy (strawberry17):

So do u think its the second one?

OpenStudy (tkhunny):

If you take out the down payment, you are even closer to something.

OpenStudy (strawberry17):

okay, thanks! It looks like it is the second one :)

OpenStudy (tkhunny):

Make sure you show the full solution when it is revealed.

OpenStudy (strawberry17):

1429.21 - 375 = $1054.21 >>>> close to $1,098.96

OpenStudy (strawberry17):

Unfortunately, it was the third answer.

OpenStudy (tkhunny):

I'd love to see it motivated.

OpenStudy (strawberry17):

It did not show how to solve for the right answer.